# Decoherence of electron spin qubit during transfer between two semiconductor quantum dots at low magnetic fields

### v2

@@ Abstract
Electron shuttling is one of the currently pursued avenues towards the scalability of semiconductor quantum dot-based spin qubits. We theoretically analyze the dephasing of a spin qubit adiabatically transferred between two tunnel-coupled quantum dots. We focus on the regime where the Zeeman splitting is lower than the tunnel coupling, at which interdot tunneling with spin flip is absent, and analyze the sources of errors in spin-coherent electron transfer for Si- and GaAs-based quantum dots. Apart from the obvious effect of fluctuations in spin splitting in each dot (e.g., due to nuclear Overhauser fields) leading to finite $T_{2}^{*}$ of the stationary spin qubit, we consider effects activated by detuning sweeps aimed at adiabatic qubit transfer between the dots: failure of charge transfer caused by charge noise and phonons, spin relaxation due to enhancement of spin-orbit mixing of levels, and spin dephasing caused by low- and high-frequency noise coupling to the electron’s charge in the presence of differences in Zeeman splittings between the two dots. Our results indicate that achieving coherent transfer of electron spin in a $10\,\mu$m long dot array necessitates a large and uniform tunnel coupling, with a typical value of $2t_{c}\gtrsim 60\,\mu$eV.
@@ I Introduction
Coherent coupling of qubit registers separated by a few microns is necessary for the development of scalable quantum computers based on semiconductor quantum dots [1, 2, 3]. Moving a spin qubit across such a distance, i.e., coherent shuttling of a single electron or hole spin, is one of the possible solutions. Such shuttling can be realized by placing the electron inside a moving potential generated using surface acoustic waves [4, 5] or metallic gates [6, 7, 8, 9, 10]. The latter approach has recently allowed for long-distance charge shuttling [8] and coherent spin shuttling on distances of a few hundred nanometers [9, 10] in Si/SiGe structures.
@@
Another approach, possible when one-dimensional chains of many tunnel-coupled quantum dots are connecting the registers, relies on sequential adiabatic transfer between neighboring quantum dots [11, 12, 13, 14, 15]: the detuning between two tunnel-coupled quantum dots is changed slowly enough for the electron to move adiabatically from one dot to another, see Fig. 1(a) and (b). This method of charge transfer was realized up to four dots in GaAs, in which, however, significant dephasing was reported [16]. More recently, in Si/SiGe quantum dots, a successful transfer of charge across 8 dots [11] was shown, followed by the transfer of a spin eigenstate across a few dots [14], and the transfer of a qubit in a superposition state between the two [13] and three [15] quantum dots.
@@
In this paper, we analyze the sources of error in coherent spin qubit transfer between two tunnel-coupled quantum dots, focusing on Si- and GaAs-based systems. We consider the spin relaxation and dephasing caused by electron-phonon interaction, high- and low-frequency charge noise, and hyperfine interaction with nuclei, with special attention devoted to mechanisms that are activated by driving the systems through an anticrossing of energy levels corresponding to electron states localized in the two dots. We work in the regime of low magnetic fields, defined by requiring the qubit Zeeman splitting, $E_{Z}$, to be smaller than the gap at the tunneling-induced anticrossing, $E_{Z}<2t_{c}$, so that there are only two anticrossings (one per spin direction) in the spectrum, see Fig. 1(b). For higher $E_{Z}$, additional anticrossings between states with opposite spin appear due to spin-orbit interaction, making the spin-flip tunneling coupling finite [17, 18, 19, 20, 21, 22]. As it is very hard to achieve perfectly adiabatic (i.e., without spin-flip) qubit transfer through these anticrossings, interference effects associated with traversing multiple anticrossings are expected to strongly affect both the probability of charge transfer [23] and the spin coherence of the transferred qubit [20]. Optimizing the driving of the two-dot system can counteract these effects [20], but performing such an optimization for a chain of many QDs is incompatible with the scalability of such an approach to long-distance quantum information transfer [7].
@@
Figure 1: Schematic illustration of the paper. a) The mechanics of dot-to-dot transition, with the relevant energy scales. The arrows indicate the order of the transition from negative detuning, through avoided crossing at $\epsilon\!=\!0$ up to the point where $\epsilon\gg 2t_{c}$. b) Energy spectrum of the four-level dot-to-dot transition. Orbital ground states are denoted using solid, and orbital excited states are denoted using dashed line for spin-up (red) and spin-down (blue). b) A schematic illustration of the considered mechanisms contributing to coherent transfer error $Q$: (1) the Landau-Zener mechanism of leaving the electron behind in the left dot due to finite sweep rate $v$, (2) charge transfer error, the process of excitation near the anticrossing due to interaction with environmental noise, followed by relaxation (inelastic tunneling) at large $\epsilon$, (3) spin relaxation enhanced by spin-orbital mixing of states near the anticrossing, (4) spin dephasing due to process (2) activated by difference of Zeeman splittings in the two dots, $\Delta E_{Z}$, and (5) dephasing caused by low-frequency noise in Zeeman splitting in each dot, and noise in $\epsilon$ when $\Delta E_{Z}$ is non-zero. c) Scaling of these contributions to error $Q$ with the detuning sweep rate $v$, shown in log-log scale. At $v$, total error $Q$ is dominated by the infidelity of charge transfer resulting from the interplay between the Landau-Zener process (orange line and Eq. 9) and incoherent charge transition to higher energy state, followed by possible relaxation (blue line and Eq. 20). At low $v$ the error is dominated by the spin relaxation (red line and Eq. (22)), and loss of coherence between spin eigenstates due to low-frequency noise (pink line and Eqs (29) and (30)), or random time of inelastic tunneling from excited to the ground level (gray line and Eq. (24)). In the regime of small error $Q\ll 1$ the total error (green line in c)) is a sum of all contributions.
@@
We thus focus on a more robust case of low magnetic fields, while taking into account the influence of high- and low-frequency noises (phonons and charge noise) on the dynamics of the transferred qubit. The fact that some of the error mechanisms discussed here become suppressed with lowering of the $B$ field provides additional motivation for focusing on low $B$. For analogous reasons, we assume that the valley splitting in each of the Si-based quantum dots, $E_{VS}$, is sufficiently large [24], i.e., fulfills $E_{VS}>2t_{c}>E_{Z}$, so that the transfer of an electron initialized in the valley ground state can be treated, to a very good approximation, as if the excited valley state was absent.
@@
Let us stress that since the long-term motivation for the research presented here is achieving coherent qubit transfer across approximately 100 quantum dots, we use rather moderate values of tunnel coupling, $1002t_{c}<$ 100 $\mu$eV. For larger values, coherent dot-to-dot transfer was shown experimentally [13, 12], but having such strong and uniform couplings might be hard to maintain for all pairs of dots in a longer array of them [7].
@@
We analyze in detail five different mechanisms of errors in coherent spin transfer, which are depicted in Fig. 1(b) and (c). Besides the Landau-Zener mechanism (1), and the charge transfer error due to interaction with the environment [23, 25] (2), we discuss mechanisms activated by the finite coupling between orbital and spin degrees of freedom, in the form of spin relaxation (3) [26] and spin dephasing (4, 5). The latter is activated by a nonzero difference in Zeeman spin splitting between the dots, $\Delta E_{Z}$, due to the gradient of the magnetic field, or differences in dot $g$-factors, the presence of which introduces dephasing when time spent in each of the dots becomes random [27, 7]. As we show, this process can take place as a result of inelastic transfer to an excited state, followed by relaxation (mechanism 4 in Fig. 1(b,c)), but also during fully adiabatic transfer due to the presence of low-frequency 1/f noise in detuning [28, 29] (mechanism 5). The dependence of all the contributions to the phase error of the shuttled qubit, $Q$, on the detuning sweep rate, $v$, is presented schematically in Fig. 1(c). One can see that depending on their relative importance, $Q(v)$ can be quite nontrivial: there can be two local minima, and both $Q\propto 1/v$ and $Q\propto 1/v^{2}$ scalings with lowering $v$ are possible.
@@
The paper is structured as follows: after using Sec. II to introduce the description of the four-level adiabatic transition, the adiabatic master equation used throughout the paper, and the model of DQD dots, in Sec. III, we briefly introduce the most relevant sources of coherent transfer error: charge transfer error (Sec. III.1), spin relaxation (Sec. III.2), and spin dephasing due to charge noise and fluctuating spin splittings (Sec. III.3). This is followed by Sec. IV, in which we compare the results for Si- and GaAs-based double quantum dots. The paper is concluded by Sec. V, in which we discuss the implications of the results for mid-range coherent transfer.
@@ II.1 Electron spin qubit in a double quantum dot
$\hat{V}_{\text{so}}=\frac{b_{\perp}}{2}\hat{\tau}_{z}\hat{\sigma}_{x}+\frac{\alpha_{\text{im}}}{2}\hat{\tau}_{y}\hat{\sigma}_{x}+\frac{\alpha_{r}}{2}\hat{\tau}_{y}\hat{\sigma}_{y},$ (4)
where the $b_{\perp}$ is caused by the difference of transverse magnetic fields between the QDs due to a magnetic field gradient or a g-tensor difference (synthetic spin-orbit interaction), and remaining terms are due to intrinsic spin-orbit coupling.
We assume here $2t_{c}>E_{Z}$, in which states with opposite spins are always separated in energy. The above spin-orbital coupling introduces the only a weak mixing of the eigenstates |$\ket{\psi_{sq}(t)}$ of Hamiltonian (1), such that the instantaneous eigenstates of total Hamiltonian ${\hat{H}_{0}(t)+\hat{V}_{\mathrm{so}}}$ can be approximated as:
@@
where |$\ket{\psi_{sq}(t)}$ are the eigenstates of $\hat{H}_{0}(t)$ with energy $E_{sq}(t)$ given in Eq. (2). Note that in considered regime of $2t_{c}>E_{Z}$, denominator is non-zero for any time instant. This allows us to neglect there modification of the $\hat{H}_{0}$ spectrum due to presence of $\hat{V}_{\text{so}}$. Since $\bra{\psi_{s^{\prime}q^{\prime}}}\hat{V}_{\text{so}}\ket{\psi_{sq}}\ll 2t_{c}-E_{Z}$, i.e. the contribution from $\hat{V}_{\text{so}}$ is relatively small, we use it only to dress the spin-diagonal states $|\psi_{sq}^{\prime}(t)\rangle$ in Eq. (5), for which now the prime will be skipped from now.
@@ II.2 Interdot transfer of the spin qubit
|$\ket{\Psi(0)}=\frac{|\psi_{\uparrow-}^{\prime}(0)\rangle+|\psi_{\downarrow-}^{\prime}(0)\rangle}{\sqrt{2}}\approx\ket{L}\otimes\frac{\ket{\uparrow}+\ket{\downarrow}}{\sqrt{2}}\,\,.$ (6)
The transfer of the qubit from L to R dot is driven by a linear sweep of detuning with sweep rate $v$, i.e. $\epsilon(t)=-\Delta\epsilon/2+vt$, from $t\!=\!0$ to $t\!=\!t_{f}\!=\!\Delta\epsilon/v$, where $\Delta\epsilon$ is the detuning swing that causes the interdot chagre transfer. It is important to note that the detuning sweep is symmetric with respect to $\epsilon\!=\!0$, i.e. changed from $-\Delta\epsilon/2$ to $\Delta\epsilon/2$.
We concentrate on a single figure of merit: the coherence between two lowest in energy instantaneous states at the final time, $t_{f}$, which for $2t_{c}\!<\!E_{Z}$ is simple the spin coherence of an electron in the ground orbital state localized in the $R$ dot:
$W(t)=2\langle\psi_{\uparrow-}^{\prime}(t_{f})|\hat{\varrho}(t_{f})|\psi_{\downarrow-}^{\prime}(t_{f})\rangle\,\,,$ (7)
where |$\ket{\psi_{s-}^{\prime}(t_{f})}$, with $s=\uparrow,\downarrow$ are two lowest-energy instantaneous eigenstates of $\hat{H}(t_{f})$, previously defined in Eq. (5). For $\epsilon(t_{f})\!\gg\!|t_{c}|$ they are very well approximated by |$\ket{R}\otimes\ket{s}$, i.e. the Zeeman doublet of states of qubit located in the $R$ dot. In the above equation $\hat{\rho}(t_{f})$ is the density matrix of the qubit in the double quantum dot. For a closed four-level system, it is given simply by $\hat{\rho}(t_{f})=\ket{\Psi(t_{f})}\bra{\Psi(t_{f})}$ |, where |$\ket{\Psi(t_{f})}$ is the state that evolves from |$\ket{\Psi(0)}$ from Eq. (6) under influence of the time-dependent Hamiltonian, $\hat{H}(t)$. Note that the factor of $2$ in Eq. (7) makes $W$ normalized in such a way that $W\!=\!1$ corresponds to maximal overlap of the final state with |$\ket{R}\otimes(\ket{\uparrow}+\ket{\downarrow})/\sqrt{2}$, i.e. a perfect transfer of spin qubit from one dot to another.
@@
According to this definition, failure of charge transfer (i.e. occupation of |$\ket{L}$ orbital state smaller than $1$ at the end of detuning sweep), spin relaxation, and spin dephasing, all contribute to $Q$. Note that the modulus in the above definition removes the deterministic phase acquired by the spin superposition state during the interdot transfer of the qubit.
@@
and shown as a yellow line in Fig. 1c. Note that we use unit in which $\hbar\!=\!1$.

### v3

@@ Abstract
Electron shuttling is one of the current avenues being pursued to scale semiconductor quantum dot-based spin qubits. Adiabatic spin qubit transfer along a chain of tunnel-coupled quantum dots is one of the possible schemes. In this scheme, we theoretically analyze the dephasing of a spin qubit that is adiabatically transferred between two tunnel-coupled quantum dots. We focus on the regime where the Zeeman splitting is lower than the tunnel coupling, such that interdot tunneling with spin flip is absent. We analyze the sources of errors in spin-coherent electron transfer for Si- and GaAs-based quantum dots. In addition to the obvious effect of fluctuations in spin splitting within each dot, leading to finite $T_{2}^{*}$ for the stationary spin qubit, we consider the effects activated by detuning sweeps: failure of charge transfer due to charge noise and phonons, spin relaxation due to the enhancement of spin-orbit mixing at the tunnel-induced anticrossing of states localized in the two dots, and spin dephasing caused by low- and high-frequency noise coupling to the electron’s charge. We show that the latter effect is activated by differences in Zeeman splittings between the two dots. Importantly, all the error mechanisms are more dangerous at low tunnel couplings. Our results indicate that away from micromagnets, maximizing the fidelity of coherent transfer aligns with minimizing charge transfer error that was previously considered in J. A. Krzywda and Ł. Cywiński, Phys. Rev. B 104 075439 (2021). For silicon, we suggest having tunnel coupling fulfilling $2t_{c}\gtrsim 60\,\mu$eV when one aims to coherently transfer a spin qubit across a $\sim\!10$ $\mu$m long array of $\sim\!100$ quantum dots with error less than $10^{-3}$.
@@ I Introduction
Coherent coupling of qubit registers separated by a few microns is necessary for the development of scalable quantum computers based on semiconductor quantum dots [1, 2, 3]. Moving a spin qubit across such a distance, i.e., coherent shuttling of a single electron or hole spin, is one of the possible solutions. Such shuttling can be realized by placing the electron inside a moving potential generated using surface acoustic waves [4, 5] or metallic gates [6, 7, 8, 9, 10]. The latter approach has recently allowed for long-distance charge shuttling [8] and coherent spin shuttling on distances of a few hundred nanometers [9, 10, 11] in Si/SiGe structures.
@@
Another approach, possible when one-dimensional chains of many tunnel-coupled quantum dots are connecting the registers, relies on sequential adiabatic transfer between neighboring quantum dots [12, 13, 14, 15, 16, 17, 11]: the detuning between two tunnel-coupled quantum dots is changed slowly enough for the electron to move adiabatically from one dot to another, see Fig. 1(a) and (b). This method of charge transfer was realized up to four dots in GaAs, in which, however, significant spin dephasing was reported [18]. More recently, in Si/SiGe quantum dots, a successful transfer of charge across 8 dots [12] was shown, followed by the transfer of a spin eigenstate across a few dots [15], and the transfer of a qubit in a superposition state between two [14, 13, 17, 19] and five [11] quantum dots. Coherent shuttling back and forth between two dots has also recently enabled the demonstration of shuttling-based single qubit gates [20].
@@
In this paper, we analyze the sources of error in coherent spin qubit transfer between two tunnel-coupled quantum dots, focusing on Si- and GaAs-based systems. We consider spin relaxation and dephasing caused by electron-phonon interaction, high- and low-frequency charge noise, and hyperfine interaction with nuclei, with special attention devoted to mechanisms that are activated by driving the systems through an anticrossing of energy levels corresponding to electron states localized in the two dots. We work in the regime of low magnetic fields, defined by requiring the qubit Zeeman splitting, $E_{Z}$, to be smaller than the gap at the tunneling-induced anticrossing, $E_{Z}<2t_{c}$, so that there are only two anticrossings (one per spin direction) in the spectrum, see Fig. 1(b). For higher $E_{Z}$, additional anticrossings between states with opposite spin appear due to spin-orbit interaction, making the spin-flip tunneling coupling finite [21, 22, 23, 24, 25, 26]. As it is very hard to achieve perfectly adiabatic (i.e., without spin-flip) qubit transfer through these anticrossings, interference effects associated with traversing multiple anticrossings are expected to strongly affect both the probability of charge transfer [27] and the spin coherence of the transferred qubit [24]. Optimizing the driving of the two-dot system can counteract these effects [24], but performing such an optimization for a chain of many QDs is incompatible with the scalability of such an approach to long-distance quantum information transfer [7]. We thus focus on a more robust case of low magnetic fields, while taking into account the influence of high- and low-frequency noises (phonons and charge noise) on the dynamics of the transferred qubit.
@@
Figure 1: Schematic illustration of the qualitative results of the paper. a) The mechanics of dot-to-dot transition, with the relevant energy scales. The arrows indicate the order of the transition from negative detuning, through avoided crossing at $\epsilon\!=\!0$ up to the point where $\epsilon\gg 2t_{c}$. b) Energy spectrum of the four-level dot-to-dot transition. Orbital ground states are denoted using solid, and orbital excited states are denoted using dashed line for spin-up (red) and spin-down (blue). b) A schematic illustration of the considered mechanisms contributing to coherent transfer error $Q$: (1) the Landau-Zener mechanism of leaving the electron behind in the left dot due to finite sweep rate $v$, (2) incoherent charge transfer error, the process of excitation near the anticrossing due to interaction with environmental noise, followed by relaxation (inelastic tunneling) at large $\epsilon$, (3) spin relaxation enhanced by spin-orbital mixing of states near the anticrossing, (4) spin dephasing due to process (2) activated by difference of Zeeman splittings in the two dots, $\Delta E_{Z}$, and (5) dephasing caused by low-frequency noise in Zeeman splitting in each dot, and noise in $\epsilon$ when $\Delta E_{Z}$ is non-zero. c) Scaling of these contributions to error $Q$ with the detuning sweep rate $v$, shown in log-log scale. At $v$, total error $Q$ is dominated by the infidelity of charge transfer resulting from the interplay between the Landau-Zener process (orange line and Eq. 9) and incoherent charge transition to higher energy state, followed by possible relaxation (blue line and Eq. 20). At low $v$ the error is dominated by the spin relaxation (red line and Eq. (22)), and loss of coherence between spin eigenstates due to low-frequency noise (violet line and Eqs (29) and (30)), or random time of inelastic tunneling from excited to the ground level (gray line and Eq. (24)). In the regime of small error $Q\ll 1$ the total error (green line in c)) is a sum of all contributions.
@@
The fact that some of the error mechanisms discussed here become suppressed with lowering of the $B$ field provides additional motivation for focusing on low $B$. For analogous reasons, we assume that the valley splitting, $E_{VS}$, in each of the Si-based quantum dots fulfills $E_{VS}>2t_{c}>E_{Z}$, so that the transfer of an electron initialized in the valley ground state can be treated, to a very good approximation, as if the excited valley state was absent. Note that this requires that the ground-state valleys states in each pair of QDs are in fact appreciably tunnel-coupled. Current understanding of valley coupling in presence of interface disorder indicates that this should be the case not only in the regime of large and deterministic valley coupling, but also in the regime of large disorder-induced randomness in valley coupling [28, 29, 30].
@@
Let us stress that since the long-term motivation for the research presented here is achieving coherent qubit transfer across approximately 100 quantum dots, we use rather moderate values of tunnel coupling, $1002t_{c}<$ 100 $\mu$eV. For larger values, coherent dot-to-dot transfer was shown experimentally [14, 13, 17], but having such strong couplings might be hard to maintain for all pairs of dots in a longer array of them [7].
We analyze in detail five mechanisms of errors in coherent spin transfer, which are depicted in Fig. 1(b) and (c). We build on the results of previous works [27, 31], where the charge transfer error resulting from a combination of the Landau-Zener mechanism (1) and environmental interactions (2) were considered. In those studies, only a lower bound on the coherent transfer error was provided, since transfer of charge is only necessary, but not sufficient, condition for coherent communication using spin shuttling.
@@
Here, we explicitly consider the spin degree of freedom, and discuss mechanisms activated by the finite coupling between orbital and spin degrees of freedom, such as spin relaxation [32] (3) and spin dephasing (4, 5). The latter is triggered by a nonzero difference in Zeeman spin splitting between the dots, $\Delta E_{Z}$, due to a magnetic field gradient or differences in the dot $g$-factors. The presence of these factors introduces dephasing when time spent in each dot becomes random [33, 7]. As we show, this process can occur due to inelastic transfer to an excited state followed by relaxation (mechanism 4 in Fig. 1(b,c)), but also during fully adiabatic transfer due to the presence of low-frequency 1/f noise in detuning [34, 35] (mechanism 5). The dependence of all contributions to the phase error of the shuttled qubit, $Q$, on the detuning sweep rate, $v$, is presented schematically in Fig. 1(c). One can see that, depending on the relative importance of the contributions, $Q(v)$ can exhibit nontrivial behavior: there may be two local minima, and both $Q\propto 1/v$ and $Q\propto 1/v^{2}$ scalings with decreasing $v$ are possible. It should be noted that while the charge transfer error as function of $v$ exhibited qualitative differences between GaAs and Si, being monotonic in GaAs, and exhibiting local minimum for Si, the total coherent transfer error considered in this paper looks qualitatively similar in the two materials, with very low and large $v$ leading to large errors, although the dominant error mechanism at low $v$ depends on the material and the choice of parameters.
@@
The paper is structured as follows: after using Sec. II to introduce the description of the four-level adiabatic transition and the adiabatic master equation used throughout the paper, in Sec. III, we briefly introduce the most relevant sources of coherent transfer error: charge transfer error (Sec. III.1), spin relaxation (Sec. III.2), and spin dephasing due to charge noise and fluctuating spin splittings (Sec. III.3). This is followed by Sec. IV, in which we compare the predictions for Si- and GaAs-based double quantum dots. The paper is concluded by Sec. V, in which we discuss the implications of the results for mid-range coherent transfer.
@@ II.1 Electron spin qubit in a double quantum dot
$\hat{V}_{\text{so}}=\frac{b_{\perp}}{2}\hat{\tau}_{z}\hat{\sigma}_{x}+\frac{a_{\text{im}}}{2}\hat{\tau}_{y}\hat{\sigma}_{x}+\frac{a_{r}}{2}\hat{\tau}_{y}\hat{\sigma}_{y},$ (4)
where the $b_{\perp}$ is caused by the difference of transverse magnetic fields between the QDs due to a magnetic field gradient or a g-tensor difference (synthetic spin-orbit interaction), and remaining terms $a=a_{r}+i\,a_{\text{im}}$ are due to intrinsic spin-orbit coupling. Note that in Si-based quantum dots, presence of electric field in $z$ direction, and atomic roughness at Si/SiGe or Si/SiO2 interface activates both Rashba and Dresselhaus spin-orbit terms, so the spin-orbit interaction is qualitatively the same for Si- and GaAs-based QDs [36, 37]- there is only a quantitative difference, with energy scale of this interaction being smaller in Si than in GaAs.
We assume here $2t_{c}>E_{Z}$, for which states with opposite spins are always separated in energy. The above spin-orbital coupling introduces then only a weak mixing of the eigenstates |$\ket{\psi_{sq}(t)}$ of Hamiltonian (1), such that the instantaneous eigenstates of total Hamiltonian ${\hat{H}_{0}(t)+\hat{V}_{\mathrm{so}}}$ can be approximated as:
@@
where |$\ket{\psi_{sq}(t)}$ are the eigenstates of $\hat{H}_{0}(t)$ with energy $E_{sq}(t)$ given in Eq. (2). Note that in considered regime of $2t_{c}>E_{Z}$, denominator is non-zero for any time instant. This allows us to neglect there modification of the $\hat{H}_{0}$ spectrum due to presence of $\hat{V}_{\text{so}}$. Since $\bra{\psi_{s^{\prime}q^{\prime}}}\hat{V}_{\text{so}}\ket{\psi_{sq}}\ll 2t_{c}-E_{Z}$. The contribution from $\hat{V}_{\text{so}}$ is relatively small and we use it only to dress the spin-diagonal states $|\psi_{sq}^{\prime}(t)\rangle$ in Eq. (5). From now we will drop the prime, and consider all of the states dressed.
@@ II.2 Interdot transfer of the spin qubit
|$\ket{\Psi(0)}=\frac{|\psi_{\uparrow-}(0)\rangle+|\psi_{\downarrow-}(0)\rangle}{\sqrt{2}}\approx\ket{L}\otimes\frac{\ket{\uparrow}+\ket{\downarrow}}{\sqrt{2}}\,\,.$ (6)
The transfer of the qubit from L to R dot is driven by a linear and symmetric sweep of detuning with characteristic sweep rate $v$, i.e. $\epsilon(t)=-\Delta\epsilon/2+vt$, from $t\!=\!0$ to $t\!=\!t_{f}\!=\!\Delta\epsilon/v$, where $\Delta\epsilon$ is the detuning swing that causes the interdot charge transfer.
We concentrate on a single figure of merit: the coherence between two lowest in energy instantaneous states at the final time, $t_{f}$. For $2t_{c}\!<\!E_{Z}$ it is simply the spin coherence of an electron in the ground orbital
$W(t)=2\langle\psi_{\uparrow-}(t_{f})|\hat{\varrho}(t_{f})|\psi_{\downarrow-}(t_{f})\rangle\,\,,$ (7)
where |$\ket{\psi_{s-}(t_{f})}$, with $s=\uparrow,\downarrow$, are two lowest-energy instantaneous eigenstates of $\hat{H}(t_{f})$, previously defined in Eq. (5) and $\hat{\rho}(t_{f})=\ket{\Psi(t_{f})}\bra{\Psi(t_{f})}$ | (where |$\ket{\Psi(t_{f})}$ is the state that evolves from |$\ket{\Psi(0)}$ from Eq. (6)under influence of $\hat{H}(t)$) is the density matrix of the electron spin in the double quantum dot. For $\epsilon(t_{f})\!\gg\!|t_{c}|$ the target eigenstates |$\ket{\psi_{s-}(t_{f})}$ are well approximated by |$\ket{R}\otimes\ket{s}$, i.e. the Zeeman doublet of states of qubit located in the $R$ dot.
@@
According to this definition, failure of charge transfer (i.e. occupation of |$\ket{R}$ orbital state smaller than $1$ at the end of detuning sweep), spin relaxation, and spin dephasing, all contribute to $Q$. Note that the modulus in the above definition removes the deterministic phase acquired by the spin superposition state during the interdot transfer of the qubit.
@@
and shown as a yellow line in Fig. 1c. Note that we use units in which $\hbar\!=\!1$.